ceneral decy 11. yor the set o. Kem shown im ถ. following frequecy your skrich does this change nect tmen and the new median, and
Answer #1

11. Table:

x f Cumulative frequency, cf xf
5 1 1 5
4 2 3 8
3 4 7 12
2 2 9 4
1 1 10 1
Total 10 30

Median :
M = value of (n/2)th observation

= value of (10/2)th observation = value of 5th observation

From the column of cumulative frequency cf, we find that the 5th observation is 3.

Hence, the median of the data is 3.

a) Histogram:

4 525150 5453 4 AJUanbay

b) Mean =  Cr/n 30/10= 3

c) The median will remain same i.e Median = 3

Th value of mean will change.

Mean = Cr/n = (4.5 * 1 4 * 2 3 * 4 2 * 2 1 * 1)/10 29.5/10 2.95

12. = N = 50, mu = 26

Σ.r μ * N = 26 * 50 = 1300